# Pantograph: A Machine-to-Machine Interaction Interface for Advanced Theorem Proving, High Level Reasoning, and Data Extraction in Lean 4

### v1

@@ 1 Introduction
A common challenge shared by all proof assistants is that completing proofs is tedious and requires significant manual effort and expertise. Machine learning offers one potential avenue for addressing this challenge. Indeed, recent years have seen several significant efforts dedicated to using machine learning to automatically search for proofs in proof assistants (e.g., [21], [22], [5], [6], [12], [11], [8], [9], [20]). While these efforts have produced promising results, many proofs are still beyond the reach of machine learning-based automation.
In order to continue to make progress in this area, several challenges need to be addressed. One of these challenges is the need for better interfaces between proof assistants and machine learning systems. In this paper, we introduce Pantograph,111https://github.com/lenianiva/PyPantograph an API and Read-Eval-Print Loop (REPL) for Lean 4, whose primary goal is to provide a convenient interface for training and evaluating theorem proving agents. The name “Pantograph” alludes to the process of recording a proof during proof search.222A Pantograph is a mechanism for recording the movement of a pen while drawing in order to create a copy.
The main motivation for creating Pantograph is to overcome the limitations of the interface provided by the Lean 4 Language Server Protocol (LSP), which is the standard interface provided for interactive use by a human user. Although the LSP provides nice interactive feedback for a human operator of Lean 4, it suffers from a number of problems as a machine interface. The LSP interface requires its user to keep track of positions of a cursor in text, and a machine user would be burdened with parsing messages from the LSP. The LSP interface also does not show information about metavariable coupling. Moreover, there is no straightforward way to extract tactic training data from the LSP interface or perform sketching. In contrast, Pantograph is designed from the ground up as an efficient and convenient interface for machine (and especially machine learning) agents.
@@
Unlike prior work, the user can decide to solve goals independently. This enables more powerful search algorithms such as Monte Carlo Tree Search (MCTS), which have been successful in other domains (e.g., AlphaGo and AlphaZero [17, 18]), achieving superhuman performance on complex games like Go, Chess, and Shogi. To do this, Pantograph handles metavariable coupling, which is a phenomenon that complicates tree search [13].
@@ 2.1 The Lean 4 Proof Assistant
A proof assistant is a computer program that can formulate and check formal mathematical proofs. This includes Lean 4 [15], Coq [19], Isabelle [16], Aya [3], and many others. In the language of a proof assistant, every definition, theorem, or proof is a value with a type. A value is represented by an expression. A proof of a theorem is a term whose type is the theorem. A proof assistant checks the validity of a proof of a theorem by evaluating its type and checking that it matches the statement of the theorem. It does this using a set of type deduction rules.
@@ 2.2 Tree Search
The tree structure that results from using tactics to prove theorems in Lean 4 is called an and-or tree and contains two types of nodes: goal nodes (Or), where solving any descendant suffices to solve the goal, and state nodes (And) produced by tactics, where solving all descendants is required. A full proof tree for the commutativity of OR is shown in Figure 1. When applying Monte Carlo Tree Search two theorem proving, two functions are required: the policy function decides which node (i.e., goal) to explore next, and the tactic function decides which tactic to use on that goal.
The use of machine learning to automatically discharge proof goals in proof assistants is an exciting and highly active area of research. The main motivation for creating Pantograph is to create an interface that can easily be used by machine learning systems aiming to exploit this incremental tree structure to conduct mathematical reasoning.
@@ 3 Related Work
The closest related work is LeanDojo [23], which provides a Python interface for machine interaction with Lean 4. A user can use this interface to execute tactics on a current proof state in a Lean 4 process. LeanDojo can be used to save and resume proof states by storing and recalling the return value of the Dojo.run_tac interface call. LeanDojo can also run commands such as #eval and #check and run tracing over an existing proof.
@@
Like LeanDojo, Pantograph can extract information from existing proofs, including training data based on triples of goal states, tactics, and post-tactic goal states. It can also extract comment data and arbitirary expressions, both of which may be useful as additional training data. However, LeanDojo’s data extraction and proof execution units are essentially separate, which makes it impossible to extract an incomplete proof and resume from it. Pantograph supports this feature.
@@
CoqGym [22] is similar to LeanDojo but for the Coq theorem prover instead of Lean 4. CoqGym stores proofs in a tree structure and allows the user agent to execute tactics. Optionally, CoqGym can serialize proof terms into S-expressions. These same features are supported by Pantograph, but for Lean 4.
@@ 4.1 Expressions and Tactics
Human-written proofs in Lean 4 (e.g., in Mathlib) are often a mixture between expressions and tactics. As mentioned above, tactics are used to reduce a goal to one or more new subgoals. To see how expressions can be used, consider the following example.
In this example, the required witness for is directly constructed as the expression . Pantograph supports seamlessly switching between expression-based and tactic-based proof by providing a custom expr tactic. This tactic takes an expression (potentially with holes ) and assigns it to the current goal. The holes then become the new goals.
@@ 4.3 Metavariable Coupling
Figure 5 gives a full example of the above proof,
conducted with automatic mode off.
The application of the transitivity tactic creates a proof state with three
goals. Using the exact 3 tactic on the
@@ 4.6 Drafting
Pantograph supports drafting in two ways. The first is via the have tactic. have introduces a lemma or intermediate claim and creates a new goal corresponding to the lemma.
@@
The user can then execute tactics on these goals as they see fit. This feature is appealing for machine learning agents, since it allows an agent (e.g., a Generative AI agent like an LLM) to effectively draft the next step of the proof without having to dive into details about its execution.
@@ 5 Evaluation
As an illustration and evaluation of Pantograph’s capabilities, we recreated the DSP experiment from [8] mentioned in Section 3. We used both the GPT-4o [1] and the GPT-o1-preview [2] language models, with parameters given in Table 1 and ran on the theorem proving evaluation benchmark MiniF2F [24]. Each individual experiment works as follows. The language model is given the formal theorem statement from MiniF2F and is asked to generate a natural language proof. Next, the same model is provided with this natural language proof and asked to generate one or three formal proof sketches in Lean 4 (GPT-o1-preview is only asked to generate a single proof sketch, as it does not yet support multiple sketches). These sketches may contain the sorry keyword. The sketches are then fed into Pantograph’s sorry-extraction command and turned into goals, which we try to solve one by one. To attempt to solve the goals, we use the following Lean 4 tactics as hammers: aesop [13], simp, and linarith (from Mathlib [14]).
@@
The results are shown Table 2. We show results on both the Validation and Test subsets of the MiniF2F benchmark set, both with one and three proof sketch(es). We report the overall success rate, and the average number of hammer invocations and the average runtime per benchmark. Our best configuration uses GPT-4o and three requested sketches. For this configuration, the DSP system out of the box successfully proved of the theorems from MiniF2F [24]. This is with no specialized training. Note that our main goal in our evaluation is to illustrate the capabilities of Pantograph. For this reason, and because they have different type systems, we do not do a comparison with the Isabelle implementation of DSP.
@@ Acknowledgements
We thank Abdalrhman Mohamed for his input on Lean 4’s expression system. We also thank Lean 4 Zulip chat users who answered our development questions. We thank Kim Morrison for the lean-training-data repository and examples of interacting with Lean 4’s front end. This work was funded in part by a gift from Amazon Web Services and by Centaur: the Stanford Center for Automated Reasoning.

### v2

@@ 1 Introduction
A common challenge shared by all proof assistants is that completing proofs is tedious and requires manual effort and expertise. Machine learning offers one potential avenue for addressing this challenge. Indeed, recent years have seen several major efforts dedicated to using machine learning to automatically search for proofs in proof assistants (e.g., [21], [22], [5], [6], [12], [11], [8], [9], [20]). While these efforts have produced promising results, many proofs are still beyond the reach of machine learning-based automation.
In order to continue to make progress in this area, several challenges need to be addressed. One of these challenges is the need for better interfaces between proof assistants and machine learning systems. In this paper, we introduce Pantograph,111https://github.com/stanford-centaur/PyPantograph an API and Read-Eval-Print Loop (REPL) for Lean 4, whose primary goal is to provide a convenient interface for training and evaluating theorem proving agents. The name “Pantograph” alludes to the process of recording a proof during proof search.222A Pantograph is a mechanism for recording the movement of a pen while drawing in order to create a copy.
The main motivation for creating Pantograph is to overcome the limitations of the interface provided by the Lean 4 Language Server Protocol (LSP), which is the standard interface provided for interactive use by a human user. Although the LSP provides interactive feedback for a human operator of Lean 4, it suffers from a number of problems as a machine interface. The LSP interface requires its user to keep track of positions of a cursor in text, and a machine user would be burdened with tracking these redundant data. Moreover, there is no straightforward way to extract tactic training data from the LSP interface or sketch out a proof to be finished by automation tactics. In contrast, Pantograph is designed from the ground up as an efficient and convenient interface for machine (and especially machine learning) agents.
@@
Unlike prior work, the user can decide to solve goals independently. This enables more powerful search algorithms such as Monte Carlo Tree Search (MCTS), which have been successful in other domains (e.g., AlphaGo and AlphaZero [17, 18]), achieving superhuman performance on complex games like Go, Chess, and Shogi.333Although these board games are not equally difficult, the state-of-the-art algorithms for these board games all involve MCTS. To do this, Pantograph handles metavariable coupling, which is a phenomenon that complicates tree search [13].
@@
As additional evidence of its usefulness, before this paper was even published, research groups in both academia and industry were already using Pantograph for machine-assisted theorem proving.
@@ 2.1 The Lean 4 Proof Assistant
A proof assistant is a computer program that can formulate and check formal mathematical proofs. This includes Lean 4 [15], Coq [19], Isabelle [16], Aya [3], and many others. These programs operate by formulating mathematics as expressions and checking the validity of the expressions via type-theoretic rules. Proof assistants may differ in a number of ways, including their syntax and the underlying variant of type theory they use. In the language of a proof assistant, every definition, theorem, or proof is a value with a type. A value is represented by an expression. A proof of a theorem is a term whose type is the theorem. A proof assistant checks the validity of a proof of a theorem by evaluating its type and checking that it matches the statement of the theorem. It does this using a set of type deduction rules.
@@ 2.2 Tree Search
The tree structure that results from using tactics to prove theorems in Lean 4 is called an and-or tree and contains two types of nodes: goals (Or), where solving at least one descendant suffices to solve the goal, and goal states (And) produced by tactics, where solving all descendants is required. A Lean 4 proof begins with a single goal. A full proof tree for the commutativity of OR is shown in Figure 1. When applying Monte Carlo Tree Search two theorem proving, two functions are required: the policy function decides which node (i.e., goal) to explore next, and the tactic function decides which tactic to use on that goal.
The main motivation for creating Pantograph is to create an interface that can easily be used by machine learning systems aiming to exploit this incremental tree structure to conduct mathematical reasoning. Potential applications of Pantograph include automatic verified program generation, rigorous reasoning for language models, and autoformalization of mathematics results.
@@ 3 Related Work
The closest related work is LeanDojo [23], which provides a Python interface for machine interaction with Lean 4. A user can use this interface to execute tactics on a current proof state in a Lean 4 process. LeanDojo can be used to save and resume proof states. LeanDojo can also run commands such as #eval and #check and extract goal-tactic pairs over an existing proof.
@@
Like LeanDojo, Pantograph can extract information from existing proofs, including training data based on triples of goal states, tactics, and post-tactic goal states. It can also extract comment data and arbitirary expressions, both of which may be useful as additional training data. However, LeanDojo’s data extraction and proof execution units are essentially separate, which makes it impossible to extract an incomplete proof and resume from it, whereas Pantograph supports this use case.
@@
CoqGym [22] is similar to LeanDojo but for the Coq theorem prover instead of Lean 4. CoqGym stores proofs in a tree structure and allows the user agent to execute tactics. Optionally, CoqGym can serialize proof terms into S-expressions. These same features are supported by Pantograph, but for Lean 4. Pantograph also has some features unsupported by CoqGym, such as the ability to implement draft-sketch-prove and to handle metavariable coupling.
@@ 4.1 Expressions and Tactics
Pantograph enables AI agents to use the same tactics as a human operator can while interacting with Lean. Human-written proofs in Lean 4 (e.g., in Mathlib) are often a mixture between expressions and tactics. As mentioned above, tactics are used to reduce a goal to one or more new subgoals. To see how expressions can be used, consider the following example.
In this example, the required witness for is directly constructed as the expression . Pantograph supports seamlessly switching between expression-based and tactic-based proof by providing a custom expr tactic. This tactic takes an expression and assigns it to the current goal. The holes then become the new goals.
There are 3 views of a proof:
1.
Presentation View: A proof written for presentation and verification. Coupling does not exist. It may contain values with puzzling origins such as complex bounds that are not apparent before reading the entire proof.
2.
Search View: A proof viewed as the trajectory of a proof search agent traverses through while finding the proof. It may contain backtracking, coupling, and goal selection.
3.
Kernel View: A proof viewed as a set of metavariables.
Pantograph enables agents to operate in the search view and handles proofs internally in the kernel view.
@@
If a tactic cannot execute for some reason, Pantograph outputs an error message corresponding to what a human operator would see during interaction with Lean’s LSP.
@@ 4.3 Metavariable Coupling
Figure 5 gives a full example of the above proof, conducted with automatic mode off. The application of the transitivity tactic creates a proof state with three goals. Using the exact 3 tactic on the goal results in a solved proof state. Applying goal.continue then brings goals and back into the proof state, where they are no longer coupled. Each can be discharged with an additional tactic such as decide.
@@ 4.6 Drafting
Pantograph supports drafting in two ways. The first is via the have tactic. This tactic introduces a lemma or intermediate claim and creates a new goal corresponding to the lemma.
@@
The user can then execute tactics on these goals as they see fit. This feature is appealing for machine learning agents, since it allows an agent (e.g., a Generative AI agent like an LLM) to effectively draft the next step of the proof without having to dive into details about its execution. If a sketch contains type errors that cannot be rectified, Pantograph forwards the error generated by the Lean 4 Kernel to the user.
4.7 Limitations
Pantograph is limited by the functionalities available in Lean. For example, if a tactic has a bug and discards a metavariable, Pantograph cannot catch the issue until the end of the proof. Anything that cannot be expressed in Lean’s type system (e.g., Homotopy Type Theory (HoTT)) also cannot be expressed in Pantograph. Due to the tight coupling of Pantograph with Lean’s internals, non-trivial engineering effort is required to update Pantograph when Lean undergoes a major version change.
Moreover, due to the user-defined nature of tactics, distributing computation via pickling of objects in Pantograph is not trivial. For example, if two branches of a proof executing on two different machines are concluded, Pantograph does not handle the algorithmically difficult problem of uniting the two branches.
@@ 5 Evaluation
In this section, we demonstrate that Pantograph has the necessary power to implement a Draft-Sketch-Prove (DSP) solver based on GPT-4o (which, notably, is not tuned for the task) and Aesop [8] (mentioned in Section 3). DSP works as follows: A formal mathematical problem is fed into a language model instance. The language model outputs a proof skeleton outlining the major steps of the formal proof. Then a proof automation tool, which does not use machine learning, fills in the holes.
We used both the GPT-4o [1] and the GPT-o1-preview [2] language models, with parameters given in Table 1, and ran on the theorem proving evaluation benchmark MiniF2F [24]. Each individual experiment works as follows. The language model is given the formal theorem statement from MiniF2F and is asked to generate a natural language proof. Next, the same model is provided with this natural language proof and asked to generate one or three formal proof sketches in Lean 4 (GPT-o1-preview is only asked to generate a single proof sketch, as it does not yet support multiple sketches). These sketches may contain the sorry keyword. The sketches are then fed into Pantograph’s sorry-extraction command and turned into goals, which we try to solve one by one. To attempt to solve the goals, we use the following Lean 4 tactics as hammers: aesop [13], simp, and linarith (from Mathlib [14]).
@@
Having built the DSP solver, the results are shown Table 2. They are not state of the art but are better than expected since we conducted no training. The drafting feature of Pantograph made this easy to implement. We show results on both the Validation and Test subsets of the MiniF2F benchmark set, both with one and three proof sketch(es). We report the overall success rate, and the average number of hammer invocations and the average runtime per benchmark. Our best configuration uses GPT-4o and three requested sketches. For this configuration, the DSP system out of the box successfully proved of the theorems from MiniF2F [24]. Since Lean and Isabelle have different type systems, we do not do a comparison with the Isabelle implementation of DSP.
@@ 6 Conclusion
Our evaluation also demonstrates one way that formal tools like Lean can be used to address potential harm from Language Models such as the one used in the evaluation section. Language models, though powerful, still face the problem of hallucination and generation of illogical results. These can be mitigated by applying formal techniques to the results produced of language models. The draft-sketch-prove experiment is an instance of this general idea, where proof automation formally checks the potentially incorrect result generated by an LLM. In the future, Pantograph could be used for other hybrid reasoning approaches combining generative AI and formal reasoning.
@@ Acknowledgements
We thank Abdalrhman Mohamed for his input on Lean 4’s expression system. We also thank Lean 4 Zulip chat users who answered our development questions. We thank Dr. Kim Morrison for the lean-training-data repository and examples of interacting with Lean 4’s front end. We thank Dr. David Dill for his feedback in proofreading the paper. This work was funded in part by a gift from Amazon Web Services and by Centaur: the Stanford Center for Automated Reasoning.